The calorie count of a serving of food can be computed based on its composition of​ carbohydrate, fat, and protein. The calorie count C for a serving of food can be computed using the formula C=4h+9f+4p​, where h is the number of grams of carbohydrate contained in the​ serving, f is the number of grams of fat contained in the​ serving, and p is the number of grams of protein contained in the serving.

A serving of raisins contains 114 calories and 25 grams of carbohydrate. If raisins are a​ fat-free food, how much protein is provided by this serving of​ raisins?


1
Expert's answer
2022-01-22T15:16:34-0500

Solution:

Given, "C=4h+9f+4p" ...(i)

"C=114, h=25,f=0,p=?"

Putting these values in (i), we get,

"114=4(25)+9(0)+4p
\\ \Rightarrow 114=100+0+4p
\\\Rightarrow 14=4p
\\ \Rightarrow p=\dfrac{14}4
\\ \Rightarrow p=3.5\ g"